OpenStudy (anonymous):

Find the equation of the tangent line to the curve y=x√x at the point (4,8).

OpenStudy (ash2326):

@jalotaibi Can you find the derivative of \(y=x\sqrt x\) ?

OpenStudy (anonymous):

√x +(x^1/2)/2 ?

OpenStudy (ash2326):

Yeah you're correct, We know that the slope of tangent drawn at a point say (a, b) is the\(\large \frac{dy}{dx}_{(a, b)}\) of the function . so substitute x=4 and y=8 in dy/dx to find the slope of the tangent.

OpenStudy (campbell_st):

can I suggest the use of index laws for this problem before anything else \[x^a \times a^b = x^{a + b}\] so \[x \times x^{\frac{1}{2}} = x^{\frac{3}{2}}\] it just eliminates the need for the product rule..

OpenStudy (ash2326):

@jalotaibi Did you understand? Can you solve the rest of the problem?

OpenStudy (anonymous):

so i put 4 and 8 instead of x ? and then solve it ?

OpenStudy (ash2326):

put x=4, we don't have a y in the equation, so we can't put y=8 anywhere find dy/ dx value, it's the slope of your tangent

OpenStudy (anonymous):

if i put 4 the answer is 2.5

OpenStudy (anonymous):

that's the slope?

OpenStudy (ash2326):

Check again, it's not 2.5

OpenStudy (anonymous):

3

OpenStudy (ash2326):

Now you have slope of the tangent as 3 and it passes through the point (4, 8) Can you write its equation?

OpenStudy (anonymous):

i get it the answer is 3x-4 ?